1. Find an infinite series series of positive terms which sums to 3
  2. Find the exact value of 1-2+4/2!-8/3!+16/4!-31/5!+64/6!-128/7!..+(-2)n/n!
1
Expert's answer
2021-11-22T05:48:36-0500

"(1)\varSigma_{n=1}^\infin ar^{n-1}=3\\
 \\−1<r<1 \ is
\\ S=\frac{{a}_{1}}{1-r}"


where "a=first \ term \ \&\ r=common\ ratio"


"S=3"


"3=\frac{a}{1-r}"


therefore

"r=\frac{3-a}{3}"


Apply the condition for convergence to determine possible values of a

"For\ a\ series\ to\ converge: −1<r<1"


"\u22121<r<1"


"−1<\frac{3-a}{3}<1"


"\u22123<3-a<3"

"-6<\u2212a<0"

"0<a<6"


"final\ answer\\

For\ the\ series\ to\ converge, 0<a<6\ 
and −1<r<1."






(2) "1-2+\frac{4}{2!}-\frac{8}{3!}+\frac{16}{4!}-\frac{32}{5!}+\frac{64}{6!}-\frac{128}{7!}+--++\frac{-2^n}{n!}+--"


"=e^{-2}=0.135"


"=0.135"